OpenStudy (anonymous):

A system of equations is shown below: 5x - 5y = 10 3x - 2y = 2 Part A: Create an equivalent system of equations by replacing one equation with the sum of that equation and a multiple of the other. Show the steps to do this. Part B: Show that the equivalent system has the same solution as the original system of equations.

OpenStudy (anonymous):

@phi @Hero

hero (hero):

multiply the first equation by 1/2, then add the result to the second equation.

OpenStudy (anonymous):

Should i multiply 1/2 by 10 Then add the five to 2?

hero (hero):

Multiply the first equation by 1/2. In other words, divide both sides of the first equation by 2

OpenStudy (anonymous):

ok i think i get it

OpenStudy (anonymous):

but that just part A correct

OpenStudy (anonymous):

\[2.5x-2.5y=10 + 3x - 2y= 2\]

OpenStudy (anonymous):

that would equal 5.5x-4.5y=12

OpenStudy (anonymous):

er no the 10 should be a 5 so 5.5x+4.5y=7

hero (hero):

Actually, what I meant to say was multiply both sides of the first equation by 1/5 You shouldn't end up with decimals if you do that.

hero (hero):

You should try to avoid decimals at all costs.

OpenStudy (anonymous):

1x-1y=2 +3x-2y=2

OpenStudy (anonymous):

4x-3y=4

OpenStudy (anonymous):

right?

hero (hero):

Yes, correct. So what would be the equivalent system now?

OpenStudy (anonymous):

4x-3y=4?

hero (hero):

That's not a system. That's just an equation.